# LUMINA-Net: Low-light Upgrade through Multi-stage Illumination and Noise Adaptation Network for Image Enhancement

### v1

@@ Abstract
Low-light image enhancement (LLIE) is a crucial task in computer vision aimed to enhance the visual fidelity of images captured under low-illumination conditions. Conventional methods frequently struggle to mitigate pervasive shortcomings such as noise, over-exposure, and color distortion thereby precipitating a pronounced degradation in image quality. To address these challenges, we propose LUMINA-Net an advanced deep learning framework designed specifically by integrating multi-stage illumination and reflectance modules. First, the illumination module intelligently adjusts brightness and contrast levels while meticulously preserving intricate textural details. Second, the reflectance module incorporates a noise reduction mechanism that leverages spatial attention and channel-wise feature refinement to mitigate noise contamination. Through a comprehensive suite of experiments conducted on LOL and SICE datasets using PSNR, SSIM and LPIPS metrics, surpassing state-of-the-art methodologies and showcasing its efficacy in low-light image enhancement.
Index Terms:
LUMINA-Net, low light image enhancement, deep learning, reflectance refinement, noise adaptation, over-exposure, PSNR, SSIM, LPIPS
@@ I Introduction
Low-Light Image Enhancement (LLIE) has emerged as a vital component in various image-based applications, including surveillance [1, 2], autonomous vehicles [3, 4], medical imaging [5, 6], and consumer electronics [7, 8, 9], where high-fidelity images are paramount. Traditional approaches to LLIE have primarily relied on two well-established methods, histogram-based and retinex-based techniques. Histogram-based LLIE techniques analyze and modify pixel intensity distributions to adjust contrast and brightness, with Histogram Equalization (HE) being a widely used method [10, 11, 12]. Retinex-based LLIE techniques separate images into reflectance and illumination components, adjusting the latter to enhance contrast and visibility while preserving natural colors and textures [13, 14, 15, 16].
@@
The comparative analysis in Figure 1 highlights the inherent limitations of these approaches. Specifically, it reveals that overexposure remains a critical challenge, significantly degrading the quality of the reconstructed images. As depicted in Figure 1, reflectance images processed by existing methods often suffer from pronounced artifacts caused by overexposure, leading to imbalanced color distribution and loss of texture detail. This inadequacy in handling under-exposed and over-exposed regions results in outputs marred by color distortion and diminished visual fidelity.
These challenges emphasize the pressing need for more advanced and adaptive techniques that can robustly address a wide range of exposure conditions while minimizing noise and mitigating overexposure. By addressing these gaps, future methodologies can achieve superior image quality, preserving both texture and color accuracy under challenging lighting scenarios.
To address the complexities of low-light image enhancement, we propose LUMINA-Net, a novel deep learning framework that leverages paired low-light images with varying exposures. Building upon the Retinex theory, LUMINA-Net decomposes low-light images into illumination and reflectance components. A key innovation of LUMINA-Net is the integration of a Channel-Guidance (CG) Module, which employs spatial and channel attention mechanisms to refine feature extraction. This is complemented by a Color Enhancement (CE) Module that ensures natural color reproduction and an Over-Exposure Correction (OEC) Module that dynamically adjusts overexposed regions, significantly reducing artifacts and improving the visual balance of the enhanced images. Additionally, self-attention-enriched skip connections are employed to preserve global coherence and fine-grain structural details, ensuring that the enhanced outputs retain their overall context and texture fidelity.
We conducted extensive experiments on the LOL and SICE datasets to evaluate our proposed retinex-based approach for low-light image enhancement. By focusing on underexposed, well-aligned image pairs, we aim to improve visual quality and system performance in challenging lighting conditions. Leveraging deep learning and image processing techniques, our goal is to advance the state-of-the-art in low-light image enhancement and unlock the full potential of vision-based systems in low-light environments.
The contributions of this paper include,
  • We introduce LUMINA-Net, a retinex-based approach for enhancing image quality. Our model utilizes paired images and incorporates a Channel-Guidance (CG) Module with spatial and channel attention for refined feature extraction, along with a Color Enhancement (CE) Module to ensure natural color balance.
  • We propose an Over-Exposure Correction (OEC) Module to dynamically adjust over-exposed regions, incorporating a self-attention-enriched skip connection for enhanced structural context and minimal artifacts.
  • We conducted extensive experiments on the LOL and SICE datasets to demonstrate the high efficiency and superiority of our LUMINA-Net, which can effectively suppress noise and preserve fine details in the final enhanced results.
II Related Work
Low-light image enhancement remains a significant challenge in computer vision and image processing, despite advancements in camera technology. Researchers have developed various solutions to address this issue, ranging from traditional methods to cutting-edge deep learning techniques, Retinex theory, and optimization algorithms. An enhanced Vector Wiener Filter (VWF) is introduced in [17] that leverages photon noise correlation and Fourier domain signal-to-noise ratio enhancement, resulting in remarkable image quality and super-resolution capability.
Building on these foundations, further innovations have emerged that resulted in Histogram Equalization (HE) with adaptive illumination adjustment [10], an adjustable contrast stretching technique to enhance color image contrast [18]. Additionally, researchers have explored the use of deep learning techniques, such as Generative Diffusion Prior (GDP) [19] and light-effects suppression networks [20, 21, 22], to address uneven light distribution and over-enhancement. These advancements have significantly improved low-light image enhancement, but ongoing research is still needed to overcome the remaining challenges.
Recent advancements in deep learning have transformed low-light image enhancement, with state-of-the-art network architectures pushing the limits of image restoration and quality. Researchers have proposed various innovative methods, including unsupervised image de-noising using GANs [23], zero-reference approaches for noise mitigation and image enhancement [24, 25], and networks leveraging RAW image data, Channel Guidance Net [26], Fourier-based transforms [27], UNet-based [28], Two-stage Single Image De-hazing Network (TSID Net) [29] and Adaptive Illumination Estimation Network (AIE Net) [30] architectures. Additionally, techniques such as style transfer-based data generation, teacher networks, and self-supervised approaches have been explored to address challenges like overexposure, underexposure, and noise removal, ultimately leading to significant improvements in low-light image enhancement.
Researchers have proposed various Retinex theory-based methods for low-light image enhancement, leveraging diverse datasets, multi-metric evaluations, and deep learning techniques. These methods decouple illumination and reflectance components to adjust lighting, suppress noise, and revive colors, restoring image clarity and visual fidelity. Recent approaches include combining Retinex theory with self-supervised learning [13, 31, 32], zero-shot learning-based Retinex decomposition method (ZERRINNet) [33], and deep neural networks such as Decom-Net, Denoise-Net, Relight-Net, Diff-Retinex, DICNet, machine learning, and CNNs [14, 15, 34, 35] to achieve superior image restoration and enhancement outcomes in low-light environments.
III Method
@@
The proposed LUMINA-Net is a multi-stage framework designed to enhance low-light images through a synergistic approach. Initially, the Channel-Guidance (CG) Module in the Reflectance branch mitigates noise and refines features, preserving intricate details and textures often compromised in low-light conditions. Subsequently, the Color Enhancement (CE) Module in the Illumination branch modulates brightness and contrast while safeguarding texture fidelity, preventing the introduction of artificial artifacts during illumination adjustment. Finally, the Over-Exposure Correction (OEC) Module harmonizes excessively bright areas, ensuring uniform lighting and retaining information in highlighted regions, thereby producing a more balanced and visually appealing image. The overall network framework is shown in Figure 2.
III-A Preliminary
The Retinex theory models a low-light image $I$ [36] as,
@@ III-A Preliminary
where $\circ$ denotes element-wise multiplication. Illumination $(L)$ represents the light intensity in the scene, expected to be smooth and texture-less, while reflectance $(R)$ captures the inherent properties of objects, such as textures and details. Conventional illumination and reflectance estimation techniques rely on predefined, hand-crafted priors that often do not accommodate the complexity and variability of real-world scenes and lighting conditions [37]. To address this limitation, we exploit paired low-light images, $I_{1}$ and $I_{2}$ with same reflectance $R^{\prime}$ but different illuminations $L_{1}^{\prime}$ and $L_{2}^{\prime}$.
@@
The introduction of paired low-light images in LUMINA-Net injects additional constraints and valuable information, bolstering the robustness of illumination-reflectance decomposition.
III-B Proposed Method
@@
Our proposed LUMINA-Net image enhancement technique seamlessly integrates three synergistic components, the Channel-Guided (CG) Module for adaptive illumination-reflectance decomposition, the Color Enhancement (CE) Module for vibrant color restoration and refinement, and the Over-Exposure Correction (OEC) Module for balanced brightness and detail preservation. This combination enables the transformation of low-quality images into visually stunning and detailed outputs, yielding a robust and efficient image enhancement framework.
III-B1 Channel-Spatial Guidance (CG) Module
The CG Module a pivotal role in enhancing the quality of reflectance maps generated from illumination-reflectance decomposition. This module leverages both Channel Attention and Spatial Attention mechanisms to selectively emphasize crucial features and suppress noise, resulting in refined reflectance maps. Initially, the CG Module takes two low-light input images with reflectance maps $R_{1}$ and $R_{2}$ that represents the intrinsic properties of the scene under different lighting conditions.
@@ III-B1 Channel-Spatial Guidance (CG) Module
The Channel Attention Mechanism then identifies significant feature channels within these reflectance maps. This is achieved through global average pooling (GAP), which generates channel descriptors that capture the importance of each channel. These descriptors are subsequently refined via a convolutional layer and sigmoid activation, producing channel-wise weights that ensure the most critical features are emphasized.
@@
In parallel, the Spatial Attention Mechanism targets essential spatial regions within the reflectance maps, such as edges and textures. This is accomplished by performing both global average pooling and max pooling along the channel dimension, generating spatial descriptors. These descriptors are then processed through a convolutional layer and sigmoid activation, producing a spatial weight map that highlights key regions for refinement.
The synergistic combination of Channel Attention and Spatial Attention mechanisms yields a refined reflectance map from $R_{f1}$ and $R_{f2}$. This integrated approach preserves high-frequency details providing a more accurate and cleaner representation of reflectance, robust to both low-light and normal-light conditions. The CG Module significantly leverages the strengths of both attention mechanisms to enhance the reflectance refinement, mitigating noise and highlighting crucial features to achieve superior image reconstruction outcomes that surpass those of traditional Retinex-based approaches.
III-B2 Color Enhancement (CE) Module
The Color Enhancement (CE) Module plays a crucial role in refining the illumination maps $L_{1}$ and $L_{2}$. Its primary objective is to enhance the brightness and contrast of the image, particularly in low-light regions, while ensuring natural and smooth lighting transitions.
@@ III-B2 Color Enhancement (CE) Module
To achieve this, the module employs a series of convolutional layers to improve the quality of the illumination maps. Additionally, it utilizes Adaptive Average Pooling (AAP), which produces a global descriptor for each channel. The descriptor is then processed through a fully connected layer and sigmoid activation, generating channel-wise attention weights. These weights modulate the features of the illumination maps, enhancing important features while suppressing irrelevant information.
The refined illumination maps, $L_{f1}$ and $L_{f2}$, are computed using Adaptive Average Pooling (AAP), which adaptively aggregates spatial information to enhance the quality of the illumination estimates. This process improves the accuracy and quality of the illumination estimates by leveraging attention mechanisms, which focus on the most relevant features for illumination correction. The channel-attended features help adaptively adjust the illumination maps, refining them for better performance in both low-light and normal-light conditions. The result is a more visually appealing and well-lit image, with enhanced clarity and detail, which can be applied in various lighting scenarios while preserving the scene’s natural appearance.
III-B3 Over-Exposure Correction (OEC) Module
@@
The Over-Exposure Correction (OEC) Module is designed to address the issue of overexposure by refining the combined illumination $(L_{f})$ and reflectance $(R_{f})$ maps. This process restores details in overexposed regions, ensuring a balanced exposure through an intermediate image representation, generated by element-wise multiplying the refined illumination and reflectance map. This intermediate representation is then passed through the OEC Module, which is specifically designed to handle areas of the image that are excessively bright or saturated.
The module employs residual learning along with sigmoid activation to suppress overexposure artifacts while preserving fine image details. The output of the OEC Module is the final enhanced image $I_{f}$, which retains natural brightness and structure while correcting overexposed regions and resulting in a well-balanced image with improved exposure.
III-C Loss Functions
@@ III-C2 Consistency Loss ($L_{C}$)
III-C3 Retinex Loss ($L_{R}$)
In this approach, the core concept is to decompose low-light images into two distinct components: illumination and reflectance. The goal is to estimate the reflectance map that captures the intrinsic properties of the scene and the illumination map that represents the varying lighting conditions. To achieve this, a set of constraints is applied to ensure that the decomposition is consistent and physically meaningful, resulting in a high-quality enhancement of low-light images. The loss function that facilitates this decomposition is formulated as follows:
@@ III-C3 Retinex Loss ($L_{R}$)
where $i$ denotes the input low-light image, $R_{f1}$ and $L_{f1}$ are the predicted reflectance and illumination maps, $L_{0}$ is the initial estimate of the illumination, and $\nabla L$ represents the gradient of the illumination map. $||L-L_{0}||_{2}^{2}$, forces the illumination estimate to be close to the initial illumination $L_{0}$, which is computed using the maximum values of the R, G, and B channels of the input image. This initialization serves as a prior, guiding the network toward more realistic illumination estimates. $||\nabla L||_{1}$, is a smoothness constraint that encourages the illumination map to have smooth transitions, avoiding sharp or unnatural changes in lighting. By minimizing this loss function, the network effectively separates the reflectance and illumination components, improving the overall quality and realism of the enhanced low-light image without relying on extensive handcrafted priors or ground-truth data.
III-C4 Perceptual Loss ($L_{per}$)
The $L_{per}$ measures the similarity between the predicted and ground-truth images in a high level feature space, capturing their perceptual quality. This loss retains essential visual details during image reconstruction by comparing features extracted from a pre-trained model.
$L_{per}=||(\phi(I_{f1}),\phi(I_{GT}))||_{2}^{2},$ (6)
where $I_{f1}$ is the predicted image, $I_{GT}$ is the ground-truth image, and $\phi(\cdot)$ represents the feature extraction function of the pretrained model.
III-D Combined Loss
The final loss function is a meticulously crafted weighted sum of individual losses, synergistically training the model to minimize perceptual, consistency, projection, reflectance, total variation and edge losses while preserving naturalness and spatial consistency. The model achieves a harmonious balance between visual fidelity, detail retention, and spatial coherence, ultimately enhancing its performance and generating high-quality images.
$L_{\text{All}}=w_{0}\times L_{p}+w_{1}\times L_{C}+w_{2}\times L_{R}$
$+w_{3}\times L_{per},$ (7)
where $L_{p}$ is the projection loss, $L_{C}$ is the reflectance consistency loss, $L_{R}$ is the retinex loss, $L_{pre}$ is the perceptual loss, and $w_{0},w_{1},w_{2},w_{3}$ are weights controlling the importance of each loss term.
IV Experiments
@@
FourierDiff [51] U 18.67 0.602 0.362 16.87 0.768 0.387
NeRco [52] U 22.94 0.773 0.327 17.97 0.756 0.428
LUMINA-Net (Ours) U 23.67 0.814 0.194 22.17 0.857 0.121
@@
IV-A Experimental Datasets

### v2

@@ Abstract
Low-light image enhancement (LLIE) is a crucial task in computer vision aimed at enhancing the visual fidelity of images captured under low-illumination conditions. Conventional methods frequently struggle with noise, overexposure, and color distortion, leading to significant image quality degradation. To address these challenges, we propose LUMINA-Net, an unsupervised deep learning framework that learns adaptive priors from low-light image pairs by integrating multi-stage illumination and reflectance modules. To assist the Retinex decomposition, inappropriate features in the raw image can be removed using a simple self-supervised mechanism. First, the illumination module intelligently adjusts brightness and contrast while preserving intricate textural details. Second, the reflectance module incorporates a noise reduction mechanism that leverages spatial attention and channel-wise feature refinement to mitigate noise contamination. Through extensive experiments on LOL and SICE datasets, evaluated using PSNR, SSIM, and LPIPS metrics, LUMINA-Net surpasses state-of-the-art methods, demonstrating its efficacy in low-light image enhancement.
@@ I Introduction
Low-Light Image Enhancement (LLIE) has emerged as a vital component in various image-based applications, including surveillance [1, 2], autonomous vehicles [3, 4], medical imaging [5, 6], and consumer electronics [7, 8, 9], where high-fidelity images are paramount. Traditional approaches to LLIE have primarily relied on two well-established methods, histogram-based and retinex-based techniques. Histogram-based LLIE techniques analyze and modify pixel intensity distributions to adjust contrast and brightness, with Histogram Equalization (HE) being a widely used method [10, 11, 12]. Retinex-based LLIE techniques separate images into reflectance and illumination components, adjusting the latter to enhance contrast and visibility while preserving natural colors and textures [13, 14, 15].
@@
The comparative analysis in Fig. 1 highlights the inherent limitations of these approaches. Specifically, it reveals that overexposure remains a critical challenge, significantly degrading the quality of the reconstructed images. As depicted in Fig. 1, reflectance images processed by existing methods often suffer from pronounced artifacts caused by overexposure, leading to imbalanced color distribution and loss of texture detail. This inadequacy in handling under-exposed and over-exposed regions results in outputs marred by color distortion and diminished visual fidelity.
These challenges highlight the need for advanced, adaptive methods that handle diverse exposure conditions, reduce noise, and mitigate overexposure—ultimately enhancing image quality while preserving texture and color accuracy in difficult lighting.
@@
To address the challenges of low-light image enhancement, we propose LUMINA-Net, a deep learning framework that utilizes paired low-light images with varying exposures. Inspired by Retinex theory, LUMINA-Net decomposes images into illumination and reflectance components. A key feature is the CG Module, which uses spatial and channel attention to refine feature extraction. The CE and OEC Modules ensure natural color reproduction and dynamic adjustment of overexposed regions, reducing artifacts. Additionally, self-attention-based skip connections preserve global coherence and fine details.
Despite progress in deep learning techniques like Generative Diffusion Prior (GDP) [16], unsupervised denoising [17], and noise-mitigation networks [18], overexposure and noise remain persistent challenges, often degrading image quality. Traditional methods like Retinex-based or histogram-based approaches do not fully address these issues. LUMINA-Net aims to overcome these limitations, improving both visual quality and system performance under challenging lighting conditions.
II Method
We define the problem of LLIE using paired low-light images and emphasize the need for an effective enhancement approach. LUMINA-Net employs a structured pipeline that refines illumination, restores color accuracy, and balances exposure. The network enhances visibility while preserving fine details and reducing unwanted distortions, ensuring natural-looking results. Carefully designed loss functions guide optimization, improving stability and overall image quality. Finally, we introduce the training dataset, which includes diverse lighting conditions to enhance model generalization. The overall network structure is illustrated in Fig. 2.
II-A Preliminary
According to the Retinex theory, low-light image $I$ can be decomposed into illumination L and reflectance R as,
@@ II-A Preliminary
where $\circ$ denotes element-wise multiplication. Illumination $(L)$ represents the light intensity in the scene, expected to be smooth and texture-less, while reflectance $(R)$ captures the inherent properties of objects, such as textures and details. Conventional illumination and reflectance estimation techniques rely on predefined, hand-crafted priors that often do not accommodate the complexity and variability of real-world scenes and lighting conditions [19, 20]. To address this limitation, we utilize paired low-light images, $I_{1}$ and $I_{2}$ with same reflectance $R^{\prime}$ but different illuminations $L_{1}^{\prime}$ and $L_{2}^{\prime}$.
@@
With paired low-light instances, we propose LUMINA-Net that injects additional constraints and valuable information, bolstering the robustness of illumination-reflectance decomposition.
II-B Proposed Method
@@
Our proposed LUMINA-Net image enhancement technique seamlessly integrates three synergistic components, the CG Module for adaptive illumination-reflectance decomposition, the CE Module for vibrant color restoration and refinement, and the OEC Module for balanced brightness and detail preservation. This combination enables effective low-light image enhancement by improving brightness, reducing noise, and maintaining structural fidelity across diverse scenes.
II-B1 Channel-Spatial Guidance (CG) Module
The CG Module plays a crucial role in refining reflectance maps generated from the illumination-reflectance decomposition. It employs both Channel Attention and Spatial Attention mechanisms to selectively highlight important features and suppress noise, ensuring better quality reflectance maps.
@@ II-B1 Channel-Spatial Guidance (CG) Module
Channel Attention Mechanism works by first applying global average pooling (GAP) to the reflectance maps $R_{1}$ and $R_{2}$ to generate channel descriptors. These descriptors are refined via a convolutional layer and sigmoid activation to produce channel-wise weights, which prioritize essential features in the reflectance maps [21].
@@
Spatial Attention Mechanism focuses on critical spatial regions within the reflectance maps, such as edges and textures. This is achieved through global average pooling and max pooling along the channel dimension, generating spatial descriptors. These descriptors are then processed through a convolutional layer and sigmoid activation to produce a spatial weight map that highlights important regions for further refinement [22].
By combining both attention mechanisms, the CG Module generates a more refined reflectance map from $R_{f1}$ and $R_{f2}$, improving the accuracy and detail preservation of the enhanced image while reducing noise. This results in a cleaner and more precise representation of the reflectance, outperforming traditional Retinex-based approaches.
II-B2 Color Enhancement (CE) Module
The CE Module refines the illumination maps $L_{1}$ and $L_{2}$ to improve brightness and contrast, especially in low-light regions, while maintaining natural lighting transitions.
@@ II-B2 Color Enhancement (CE) Module
It uses convolutional layers to enhance the illumination maps and Adaptive Average Pooling (AAP) to generate global channel descriptors. These descriptors are processed through a fully connected layer and sigmoid activation to produce channel-wise attention weights, emphasizing important features and suppressing irrelevant ones [23].
The refined illumination maps $L_{f1}$ and $L_{f2}$ are derived by aggregating spatial information through Adaptive Average Pooling (AAP), enhancing the illumination estimates. This process, driven by attention mechanisms, ensures more accurate and visually pleasing results in both low-light and normal-light conditions.
II-B3 Over-Exposure Correction (OEC) Module
@@
The OEC Module is designed to address the issue of overexposure by refining the combined illumination $(L_{f})$ and reflectance $(R_{f})$ maps. This process restores details in overexposed regions, ensuring a balanced exposure through an intermediate image representation, generated by element-wise multiplying the refined illumination and reflectance map. This intermediate representation is then passed through the OEC Module, which is specifically designed to handle areas of the image that are excessively bright or saturated. To compute the final enhanced image, we apply the illumination correction factor $\lambda$ as follows:
$I_{f}=L^{\lambda}\circ R,$ (3)
where $\lambda$ is the illumination correction factor, and $I_{f}$ denotes the enhanced image. This ensures a refined and well-balanced output by dynamically adjusting the illumination to prevent excessive brightness while maintaining structural details.
II-C Loss Functions
@@ II-C2 Consistency Loss ($L_{C}$)
II-C3 Retinex Loss ($L_{R}$)
This approach decomposes low-light images into illumination and reflectance components. The goal is to estimate the reflectance map, capturing intrinsic scene properties, and the illumination map, representing lighting conditions. Constraints are applied to ensure the decomposition is consistent and physically meaningful, resulting in high-quality enhancements. The loss function is formulated as follows:
@@ II-C3 Retinex Loss ($L_{R}$)
where $i$ is the input low‑light image, $R_{f1}$ and $L_{f1}$ are the predicted reflectance and illumination maps, $L_{0}$ is the initial illumination estimate, and $\nabla L$ represents the illumination gradient. The operator $\mathrm{stopgrad}(\cdot)$ acts as an identity in the forward pass but blocks gradient flow during backpropagation, keeping $L_{f1}$ fixed in the term $R_{f1}-i/\mathrm{stopgrad}(L_{f1})$. This prevents trivial decompositions by focusing training on $R_{f1}$. Finally, $\|L-L_{0}\|_{2}^{2}$ anchors the illumination map, and $\|\nabla L\|_{1}$ promotes smoothness. Minimizing this loss effectively separates the reflectance and illumination components, enhancing the low‑light image quality.
@@
NeRco [38] U 19.80 0.730 0.240 20.73 0.820 0.230
FourierDiff [39] U 21.94 0.710 0.362 18.87 0.768 0.387
LightenDiffusion [40] U 20.45 0.803 0.209 19.08 0.776 0.375
LUMINA-Net (Ours) U 23.92 0.812 0.180 22.65 0.853 0.131
@@
II-C4 Perceptual Loss ($L_{per}$)
The $L_{per}$ measures the similarity between the two predicted enhanced images ($I_{f1}$ and $I_{f2}$) in a high-level feature space, capturing their perceptual consistency. This encourages both enhanced outputs to retain visually coherent features by comparing their representations extracted from a pre-trained network.
$L_{per}=||(\phi(I_{f1}),\phi(I_{f2}))||_{2}^{2},$ (7)
where $I_{f1}$ and $I_{f2}$ are the predicted enhanced images from different augmented inputs, and $\phi(\cdot)$ represents the feature extraction function of the pretrained model.
II-D Combined Loss
The final loss is a weighted sum of individual losses, optimizing perceptual, consistency, projection, reflectance, and edge losses to balance visual fidelity, detail, and spatial coherence. The Combined Loss is formulated as follows:
$L_{\text{All}}=w_{0}\times L_{p}+w_{1}\times L_{C}+w_{2}\times L_{R}+w_{3}\times L_{per},$
where $L_{p}$ is projection loss, $L_{C}$ is reflectance consistency loss, $L_{R}$ is retinex loss, and $L_{per}$ is perceptual loss, with $w_{0},w_{1},w_{2},w_{3}$ as weights.
III Experiments
III-A Experimental Datasets